# NAI2: Learning Noise-Aware Illumination Interpolator for Unsupervised Low-Light Image Enhancement

### v3

@@ Abstract
Contemporary Low-Light Image Enhancement (LLIE) techniques have made notable advancements in preserving image details and enhancing contrast, achieving commendable results on specific datasets. Nevertheless, these approaches encounter persistent challenges in efficiently mitigating dynamic noise and accommodating diverse low-light scenarios. Insufficient constraints on complex pixel-wise mapping learning lead to overfitting to specific types of noise and artifacts associated with low-light conditions, reducing effectiveness in variable lighting scenarios. To this end, we first propose a method for estimating the noise level in low light images in a quick and accurate way. This facilitates precise denoising, prevents over-smoothing, and adapts to dynamic noise patterns. Subsequently, we devise a Learnable Illumination Interpolator (LII), which employs learnlable interpolation operations between the input and unit vector to satisfy general constraints between illumination and input. Finally, we introduce a self-regularization loss that incorporates intrinsic image properties and essential visual attributes to guide the output towards meeting human visual expectations. Comprehensive experiments validate the competitiveness of our proposed algorithm in both qualitative and quantitative assessments. Notably, our noise estimation method, with linear time complexity and suitable for various denoisers, significantly improves both denoising and enhancement performance. Benefiting from this, our approach achieves a 0.675dB PSNR improvement on the LOL dataset and 0.818dB on the MIT dataset on LLIE task, even compared to supervised methods.
@@
Image processing , low-light image enhancement , noise estimation , illumination learning
journal: Journal of  Templates
[…]
[label1] organization=School of Software Technology,addressline=Dalian University of Technology, city=Dalian, postcode=116620, state=Liaoning, country=China
@@ 1 Introduction
As a significant branch of image restoration, Low-Light Image Enhancement (LLIE) aims at improving the visual quality and details of images, to gather brighter, clearer, and more detailed images captured in low-light conditions. By improving the quality of images captured under low-light conditions, LLIE finds wide applications in various fields and downstream machine analytics, such as for consumer photography [1, 2], autonomous driving [3], and video surveillance [4]. Overall, the development of low-light enhancement technology holds significant importance in improving image quality, enhancing visibility, and increasing efficiency and safety across various domains.
@@
Unsupervised learning approaches in low-light enhancement aim to improve image quality without the need for labeled training data. These methods leverage statistical properties of low-light images or heuristic algorithms to enhance brightness, contrast, and details. To direct end-to-end learning for this mapping process in unsupervised way, many methods improve the brightness and contrast of low-light images by performing pixel-wise addition (e.g. adding a pixel-wise map to adjust the curve of brightness [5, 6]) or division(e.g. dividing learned illumination map [7, 8]). Despite efforts to uncover relationships and constraints among different quantities during the enhancement process, these methods do not explicitly impose constraints on the statistical distribution of data that aligns with human visual output. Consequently, training guidance solely from input data may result in overfitting to specific types of noise or artifacts. This overall limitation makes these methods less effective in variable and dynamic lighting conditions, where the types of degradation may vary.
@@
Figure 1: Comparison among recent state-of-the-art unsupervised methods, including EnGAN [9], ZeroDCE [5], RetinexDIP [10], RUAS [11], and SCI [8]. A scaled heatmap showcasing the absolute error (darker is better) and an illumination map are concatenated to the right part of the enhanced image for comprehensive comparison. Our method closely resembles the ground truth and exhibits a globally smooth yet structure-aware illumination map.
@@
In addition to learning the mapping from dark to bright, denoising is also a crucial step in LLIE. Because images captured under low-light conditions often exhibit noise, which can be attributed to factors such as low signal-to-noise ratio, high ISO values and dark current noise. While some methods [5, 8] implicitly suppress noise by imposing smoothness constraints on the illumination map, these constraints may not effectively improve the smoothness of the enhancement results. This limitation becomes particularly apparent when dealing with high-intensity noise, making the constraints appear less effective. On the other hand, certain techniques [7, 12] explicitly address noise issues by incorporating denoising components. However, they do not take into account the dynamic noise distribution in degraded images, leading the model overfit to specific noise and providing limited useful visual enhancement. To get better denoising results, previous research [13, 14] has demonstrated the effectiveness of integrating noise intensity as an input for enhancing non-blind denoising methods. How to estimate the noise intensity in a fast and accurate way in low-light condition to improve the performance is what we concerned.
@@ 2.1 Low-light Image Enhancement
In the realm of LLIE, researchers have delved into an assortment of network models aimed at tackling the amalgamation of degradation factors and improving representations. RetinexNet[15] introduced a semi-decoupled approach comprising a Decom-Net for decomposition and an Enhance-Net for illumination adjustment. MBLLEN[16] employed a deep learning method with multi-branch fusion to extract features from various levels and enhance low-light images through multiple subnets. GLADNet[17] dealt with LLIE via a Global illumination Aware and Detail-preserving Network. RUAS[11] employed a collaborative reference-free learning searching strategy to discover a suitable network architecture for enhancing low-light images, obviating the need for handcrafting design. Besides, numerous training strategies have been explored for LLIE to investigate the relationship between low-light and normal images. DRBN [18] presented a semi-supervised learning approach utilizing deep recursive band networks to extract coarse-to-fine band representations and reconstitute towards high-quality images. ZeroDCE [5] estimated pixel-wise and high-order curves for dynamic range adjustment of a given image using a lightweight deep network. EnGAN [9] introduced a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanism to train the network in an adversarial way without paired images. RetinexDIP [10] devised a novel ”generative” strategy for Retinex decomposition and a unified deep framework for low-light image enhancement, without external image data. SCI [8] put forward a new self-calibrated illumination learning framework to reduce computational cost and accelerate the training process.
@@ 2.2 Image Denoising
Previous denoising models, such as IRCNN [19] and DnCNN [20], have traditionally relied on supervised learning methods tailored to distinct noise levels, requiring separate networks for different degradation factors. In contrast, FFDNet [13] leverages noise estimation as input, demonstrating the capability to effectively handle a wide range of noise levels using a singular model. Building upon this paradigm, both DRUNet [14] and CBDNet [21] exhibit impressive denoising proficiency by integrating noise level inputs and employing a UNet-style denoising architecture. DRUNet directly incorporates noise levels as input for to solve multiple tasks related to image restoration, whereas CBDNet estimates noise levels using convolutional networks. The challenge arises from the dynamic nature of real-world noise, which is not solely static. This characteristic presents a significant hurdle for neural networks, hindering precise estimation through learning-based methods. Given the dynamic attributes inherent in real-world noise, the prompt and accurate estimation of noise levels becomes crucial. This dynamic quality poses a considerable challenge for both neural networks and conventional methods in accurately estimating the intensity of dynamic noise.
@@ 2.3 Noise Level Estimation
Estimating noise levels from a single image is considered an inherently challenging problem. Despite efforts spanning several decades, numerous methods have emerged for this purpose, often operating under the assumption that the processed image contains sufficient flat areas. However, this assumption is not consistently applicable, particularly in scenarios involving the processing of natural images. Pyatykh [22], Liu [23], and Chen [24] have explored the utilization of Principal Component Analysis (PCA) to model observed signals and infer the noise distribution. While these methods represent advancements, they still possess high time complexity in evaluating noise distribution, rendering them unsuitable for real-time enhancement as a pre-processing progress.
@@ 3 Preliminaries
Derived from the study of the human visual system [25], Retinex theory serves as a color constancy model with broad applications in mitigating problems associated with uneven lighting and color bias. The fundamental tenet of the Retinex theory involves the partitioning of an image into its constituent reflectance and illumination components. The Retinex theory views an image as the product of its reflectance and illumination components. The reflectance component represents the color and texture information of the object’s surface, while the illumination component represents the distribution of lighting. In LLIE, the relationship of low-light input $\mathbf{x}$, reflectence $\mathbf{s}$ and illumination $\mathbf{y}$ can be formulated as follows:
@@ 4 Proposed Method
Figure 2: Illustrations of the fundamental components for NAI2. Top: (a) Self-Calibrated Denoiser, (b) Learnable Illumination Interpolator. Bottom: (c) Denoise Module $\mathcal{N}_{d}(\mathbf{x},\bm{\omega_{d}})$, (d) Illumination Learning Module $\mathcal{M}_{i}(\mathbf{x},\bm{\omega_{i}})$, (e) Conditional Feature Modulation (CFM).
@@
Figure 3: Visualization of intermediate results (i.e., $\mathbf{x}$, $\mathbf{v}$, $\mathbf{y}$ and $\mathbf{s}$) regarding the workflow of NAI2 based on Equation3 (2).
@@ 4.1 First-Denoising Last-Enhancing
Motivated by the Retinex theory [25], we adapt a more general framework for low-light imaging that accounts for incidental compatible noise. The framework can be formulated as $\mathbf{x}=\mathbf{y}\otimes\mathbf{s}+\mathbf{v}$, where $\mathbf{x}$ denotes the low-light input corrupted by incidental noise $\mathbf{v}$, $\mathbf{s}$ represents the reflectance, $\mathbf{y}$ represents the illumination, and $\otimes$ denotes the element-wise product operation.
@@ 4.3 Learnable Illumination Interpolator
where $\mathbf{1}$ denotes the unit vector and $\bm{\alpha}$ is the learned illumination interpolation factor. As shown in (d) of Figure 2, $\mathcal{M}_{i}(\mathbf{u},\bm{\omega_{i}})$ is constructed as a network module. The interpolation is a natural structure constraint on fidelity and smooth properties of illumination and we will delve into the compelling advantages of LII with such structure constraint in Section Discussion 5.
@@ 4.4 Reference-Free Loss Function
where $\bar{\mathbf{s}}$ is the mean value of image $\mathbf{s}$ in different channels, $\eta_{t}$ and $\sigma_{t}$ are constants, and presents the mean and standard deviation value of the desired enhanced image. $\eta_{t}$ is set to $[0.485,0.456,0.406]$ and $\sigma_{t}$ is set to $[0.229,0.224,0.225]$ according to statistical properties of Imagenet [26], revealing a statistic value of image intensity. These values represent common attributes of typical images and hold practical value, being widely applied in preprocessing processes like detection and segmentation. Further discussion is provided in Section Discussion 7.
Self-Adaptable Noise Removal Loss. To preserve edge details during the denoising process, we introduce an adaptive denoising loss function that dynamically balances the weights of fidelity term and regularization term [27], thereby ensuring optimal performance across varying noise intensities encountered in unsupervised denoising. It can be formulate as
@@ 4.5 Discussion
Charm of Illumination Interpolator.   Compared to Retinex-induced LLIE methods, the proposed illumination interpolator learns a difference function from the original image to the illumination map. The interpolation between input and unit vector effectively performs edge-preserving smoothing on the input image. On one hand, the differences between pixels decrease after interpolation, resulting in an overall smoother appearance. On the other hand, the interpolation process does not disturb the relative spatial relationships between pixels, thus preserving edges and ensuring the overall structure of the image remains stable. As a result, the output maintains consistency in relative brightness and structural fidelity with the input. That’s to say, the output of LII naturally satisfies the general assumptions for illumination smoothing and structural fidelity, which are the main Note that comparison results of the Retinex-induced LLIE methods is illustrated in Figure 5. Considering the Retinex-induced model, we calculate the illumination map $\mathbf{y}$ of various methods by using $\mathbf{x}\oslash\mathbf{s}$. Obviously, the illumination map estimated by our NAI2 is smoother and retains the edge detail information of the original image.
@@ 5.1 Noise Estimation Evaluation
Liu. [23] 0.036 0.197 0.373 0.506 7.916
Chen. [24] 0.030 0.067 0.119 0.181 0.038
IVHC [28] 0.275 0.335 0.550 0.734 0.012
@@
To comprehensively evaluate the advantages of integrating estimated noise levels, we executed a series of experiments. In instances where denoising models like IRCNN [19] and DnCNN [20] traditionally do not consider noise levels as additional inputs, we adopted a methodology involving the embedding of our estimated noise into a vector. This embedded vector was seamlessly integrated into the shadow layer of these widely-used denoising models through the incorporation of multiple linear layers. Regarding CBDNet [21], a model typically reliant on neural network-based noise estimation, we replaced its original estimation module with our proposed noise estimation method. All models underwent training using the common Mean Squared Error (MSE) loss function for image restoration. Training procedures utilized a dataset comprising 596 clear low-light images sourced from the MIT dataset. Subsequently, testing involved 99 additional images, each augmented with Gaussian noise characterized by dynamically varying intensities. Importantly, during training, the noise intensity ranged randomly from 2 to 50, while during testing, it varied from 5 to 35. This random variation ensured that the denoising models remained oblivious to specific noise levels, maintaining a blind status throughout the entire process. Each individual model underwent training for a total of 4000 iterations. This comprehensive experimental setup facilitated a thorough evaluation of how our estimated noise impacted the denoising models’ performance.

### v4

# Advancing Unsupervised Low-light Image Enhancement: Noise Estimation, Illumination Interpolation, and Self-Regulation

@@ Abstract
Contemporary Low-Light Image Enhancement (LLIE) techniques have made notable advancements in preserving image details and enhancing contrast, achieving commendable results on specific datasets. Nevertheless, these approaches encounter persistent challenges in efficiently mitigating dynamic noise and accommodating diverse low-light scenarios. Insufficient constraints on complex pixel-wise mapping learning lead to overfitting to specific types of noise and artifacts associated with low-light conditions, reducing effectiveness in variable lighting scenarios. To this end, we first propose a method for estimating the noise level in low light images in a quick and accurate way. This facilitates precise denoising, prevents over-smoothing, and adapts to dynamic noise patterns. Subsequently, we devise a Learnable Illumination Interpolator (LII), which employs learnlable interpolation operations between the input and unit vector to satisfy general constraints between illumination and input. Finally, we introduce a self-regularization loss that incorporates intrinsic image properties and essential visual attributes to guide the output towards meeting human visual expectations. Comprehensive experiments validate the competitiveness of our proposed algorithm in both qualitative and quantitative assessments. Notably, our noise estimation method, with linear time complexity and suitable for various denoisers, significantly improves both denoising and enhancement performance. Benefiting from this, our approach achieves a 0.675dB PSNR improvement on the LOL dataset and 0.818dB on the MIT dataset on LLIE task, even compared to supervised methods. The source code is available at this DOI repository and the specific code for noise estimation can be found at this separate GitHub link.
@@
Image processing low-light image enhancement noise estimation illumination learning
@@ 1 Introduction
As a significant branch of image restoration, Low-Light Image Enhancement (LLIE) aims at improving the visual quality and details of images, to gather brighter, clearer, and more detailed images captured in low-light conditions. By improving the quality of images captured under low-light conditions, LLIE finds wide applications in various fields and downstream machine analytics, such as autonomous driving Driving , dark object detection Hu2022 Jiang2024 , and video surveillance DBLP:conf/mlsp/PotterGLHNW20 . Overall, the development of low-light enhancement technology holds significant importance in improving image quality, enhancing visibility, and increasing efficiency and safety across various domains.
@@
Unsupervised learning approaches in low-light enhancement aim to improve image quality without the need for labeled training data. These methods leverage statistical properties of low-light images or heuristic algorithms to enhance brightness, contrast, and details. To direct end-to-end learning for this mapping process in unsupervised way, many methods improve the brightness and contrast of low-light images by performing pixel-wise addition (e.g. adding a pixel-wise map to adjust the curve of brightness guo2020zero ) or division(e.g. dividing learned illumination map liu2021retinex ma2022toward ). Despite efforts to uncover relationships and constraints among different quantities during the enhancement process, these methods do not explicitly impose constraints on the statistical distribution of data that aligns with human visual output. Consequently, training guidance solely from input data may result in overfitting to specific types of noise or artifacts. This overall limitation makes these methods less effective in variable and dynamic lighting conditions, where the types of degradation may vary.
@@
Figure 1: Comparison among recent state-of-the-art unsupervised methods, including EnGAN jiang2021enlightengan , ZeroDCE guo2020zero , RetinexDIP zhao2021retinexdip , RUAS liu2021ruas , and SCI ma2022toward . A scaled heatmap showcasing the absolute error (darker is better) and an illumination map are concatenated to the right part of the enhanced image for comprehensive comparison. Our method closely resembles the ground truth and exhibits a globally smooth yet structure-aware illumination map.
@@
In addition to learning the mapping from dark to bright, denoising is also a crucial step in LLIE. Because images captured under low-light conditions often exhibit noise, which can be attributed to factors such as low signal-to-noise ratio, high ISO values and dark current noise. While some methods guo2020zero ma2022toward implicitly suppress noise by imposing smoothness constraints on the illumination map, these constraints may not effectively improve the smoothness of the enhancement results. This limitation becomes particularly apparent when dealing with high-intensity noise, making the constraints appear less effective. On the other hand, certain techniques liu2021retinex explicitly address noise issues by incorporating denoising components. However, they do not take into account the dynamic noise distribution in degraded images, leading the model overfit to specific noise and providing limited useful visual enhancement. To get better denoising results, previous research zhang2018ffdnet zhang2021plug has demonstrated the effectiveness of integrating noise intensity as an input for enhancing non-blind denoising methods. How to estimate the noise intensity in a fast and accurate way in low-light condition to improve the performance is what we concerned.
@@ 2.1 Low-light Image Enhancement
In the realm of LLIE, researchers have delved into an assortment of network models aimed at tackling the amalgamation of degradation factors and improving representations. RetinexNetwei2018deep introduced a semi-decoupled approach comprising a Decom-Net for decomposition and an Enhance-Net for illumination adjustment. MBLLENlv2018mbllen employed a deep learning method with multi-branch fusion to extract features from various levels and enhance low-light images through multiple subnets. GLADNetwang2018gladnet dealt with LLIE via a Global illumination Aware and Detail-preserving Network. RUASliu2021ruas employed a collaborative reference-free learning searching strategy to discover a suitable network architecture for enhancing low-light images, obviating the need for handcrafting design. Besides, numerous training strategies have been explored for LLIE to investigate the relationship between low-light and normal images. DRBN yang2020fidelity presented a semi-supervised learning approach utilizing deep recursive band networks to extract coarse-to-fine band representations and reconstitute towards high-quality images. ZeroDCE guo2020zero estimated pixel-wise and high-order curves for dynamic range adjustment of a given image using a lightweight deep network. EnGAN jiang2021enlightengan introduced a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanism to train the network in an adversarial way without paired images. RetinexDIP zhao2021retinexdip devised a novel ”generative” strategy for Retinex decomposition and a unified deep framework for low-light image enhancement, without external image data. SCI ma2022toward put forward a new self-calibrated illumination learning framework to reduce computational cost and accelerate the training process.
@@ 2.2 Image Denoising
Previous denoising models, such as IRCNN IRCNN and DnCNN DnCNN , have traditionally relied on supervised learning methods tailored to distinct noise levels, requiring separate networks for different degradation factors. In contrast, FFDNet zhang2018ffdnet leverages noise estimation as input, demonstrating the capability to effectively handle a wide range of noise levels using a singular model. Building upon this paradigm, both DRUNet zhang2021plug and CBDNet CBDNet exhibit impressive denoising proficiency by integrating noise level inputs and employing a UNet-style denoising architecture. DRUNet directly incorporates noise levels as input for to solve multiple tasks related to image restoration, whereas CBDNet estimates noise levels using convolutional networks. The challenge arises from the dynamic nature of real-world noise, which is not solely static. This characteristic presents a significant hurdle for neural networks, hindering precise estimation through learning-based methods. Given the dynamic attributes inherent in real-world noise, the prompt and accurate estimation of noise levels becomes crucial. This dynamic quality poses a considerable challenge for both neural networks and conventional methods in accurately estimating the intensity of dynamic noise.
@@ 2.3 Noise Level Estimation
Estimating noise levels from a single image is considered an inherently challenging problem. Despite efforts spanning several decades, numerous methods have emerged for this purpose, often operating under the assumption that the processed image contains sufficient flat areas. However, this assumption is not consistently applicable, particularly in scenarios involving the processing of natural images. Pyatykh NoisePCA2012TIP , Liu NoiseLevelEstimation2013TIP , and Chen StatisticalMethod2015TIP have explored the utilization of Principal Component Analysis (PCA) to model observed signals and infer the noise distribution. While these methods represent advancements, they still possess high time complexity in evaluating noise distribution, rendering them unsuitable for real-time enhancement as a pre-processing progress.
@@ 3 Preliminaries
Derived from the study of the human visual system land1977retinex , Retinex theory serves as a color constancy model with broad applications in mitigating problems associated with uneven lighting and color bias. The fundamental tenet of the Retinex theory involves the partitioning of an image into its constituent reflectance and illumination components. The Retinex theory views an image as the product of its reflectance and illumination components. The reflectance component represents the color and texture information of the object’s surface, while the illumination component represents the distribution of lighting. In LLIE, the relationship of low-light input $\mathbf{x}$, reflectence $\mathbf{s}$ and illumination $\mathbf{y}$ can be formulated as follows:
@@ 4 Proposed Method
Figure 2: Illustrations of the pipeline and fundamental components. Top: (a) Self-Calibrated Denoiser, (b) Learnable Illumination Interpolator. Bottom: (c) Denoise Module $\mathcal{N}_{d}(\mathbf{x},\bm{\omega_{d}})$, (d) Illumination Learning Module $\mathcal{M}_{i}(\mathbf{x},\bm{\omega_{i}})$, (e) Conditional Feature Modulation (CFM).
@@
Figure 3: Visualization of intermediate results (i.e., $\mathbf{x}$, $\mathbf{v}$, $\mathbf{y}$ and $\mathbf{s}$) regarding the workflow based on Equation3 (2).
@@ 4.1 First-Denoising Last-Enhancing
Motivated by the Retinex theory land1977retinex , we adapt a more general framework for low-light imaging that accounts for incidental compatible noise. The framework can be formulated as $\mathbf{x}=\mathbf{y}\otimes\mathbf{s}+\mathbf{v}$, where $\mathbf{x}$ denotes the low-light input corrupted by incidental noise $\mathbf{v}$, $\mathbf{s}$ represents the reflectance, $\mathbf{y}$ represents the illumination, and $\otimes$ denotes the element-wise product operation.
@@ 4.3 Learnable Illumination Interpolator
where $\mathbf{1}$ denotes the unit vector and $\bm{\alpha}$ is the learned illumination interpolation factor. As shown in (d) of Figure 2, $\mathcal{M}_{i}(\mathbf{u},\bm{\omega_{i}})$ is constructed as a network module. The interpolation is a natural structure constraint on fidelity and smooth properties of illumination and we will delve into the compelling advantages of LII with such structure constraint in Section Discussion 4.5.
@@ 4.4 Reference-Free Loss Function
where $\bar{\mathbf{s}}$ is the mean value of image $\mathbf{s}$ in different channels, $\eta_{t}$ and $\sigma_{t}$ are constants, and presents the mean and standard deviation value of the desired enhanced image. $\eta_{t}$ is set to $[0.485,0.456,0.406]$ and $\sigma_{t}$ is set to $[0.229,0.224,0.225]$ according to statistical properties of Imagenet deng2009imagenet , revealing a statistic value of image intensity. These values represent common attributes of typical images and hold practical value, being widely applied in preprocessing processes like detection and segmentation. Further discussion is provided in Section Discussion 7.
Self-Adaptable Noise Removal Loss. To preserve edge details during the denoising process, we introduce an adaptive denoising loss function that dynamically balances the weights of fidelity term and regularization term rudin1992nonlinear , thereby ensuring optimal performance across varying noise intensities encountered in unsupervised denoising. It can be formulate as
@@ 4.5 Discussion
Charm of Illumination Interpolator.   Compared to Retinex-induced LLIE methods, the proposed illumination interpolator learns a difference function from the original image to the illumination map. The interpolation between input and unit vector effectively performs edge-preserving smoothing on the input image. On one hand, the differences between pixels decrease after interpolation, resulting in an overall smoother appearance. On the other hand, the interpolation process does not disturb the relative spatial relationships between pixels, thus preserving edges and ensuring the overall structure of the image remains stable. As a result, the output maintains consistency in relative brightness and structural fidelity with the input. That’s to say, the output of LII naturally satisfies the general assumptions for illumination smoothing and structural fidelity, which are the main Note that comparison results of the Retinex-induced LLIE methods is illustrated in Figure 5. Considering the Retinex-induced model, we calculate the illumination map $\mathbf{y}$ of various methods by using $\mathbf{x}\oslash\mathbf{s}$. Obviously, the illumination map estimated by ours is smoother and retains the edge detail information of the original image.
@@ 5.1 Noise Estimation Evaluation
Liu. NoiseLevelEstimation2013TIP 0.036 0.197 0.373 0.506 7.916
Chen. StatisticalMethod2015TIP 0.030 0.067 0.119 0.181 0.038
IVHC IVHC2016TIP 0.275 0.335 0.550 0.734 0.012
@@
To comprehensively evaluate the advantages of integrating estimated noise levels, we executed a series of experiments. In instances where denoising models like IRCNN IRCNN and DnCNN DnCNN traditionally do not consider noise levels as additional inputs, we adopted a methodology involving the embedding of our estimated noise into a vector. This embedded vector was seamlessly integrated into the shadow layer of these widely-used denoising models through the incorporation of multiple linear layers. Regarding CBDNet CBDNet , a model typically reliant on neural network-based noise estimation, we replaced its original estimation module with our proposed noise estimation method. All models underwent training using the common Mean Squared Error (MSE) loss function for image restoration. Training procedures utilized a dataset comprising 596 clear low-light images sourced from the MIT dataset. Subsequently, testing involved 99 additional images, each augmented with Gaussian noise characterized by dynamically varying intensities. Importantly, during training, the noise intensity ranged randomly from 2 to 50, while during testing, it varied from 5 to 35. This random variation ensured that the denoising models remained oblivious to specific noise levels, maintaining a blind status throughout the entire process. Each individual model underwent training for a total of 4000 iterations. This comprehensive experimental setup facilitated a thorough evaluation of how our estimated noise impacted the denoising models’ performance.